OpenStudy (anonymous):

Given the linear equation y=-2/3x+1, find the y-coordinates of the points (-6, ), (-3, ), and (3, ). Please show all of your work.

OpenStudy (australopithecus):

so you have x = -6 x = -3 x = 3 plug and chug

OpenStudy (australopithecus):

remember a point is (x,y) you have (-6, y) (-3, y) (3, y)

OpenStudy (australopithecus):

plug the numbers in as x and solve for y

OpenStudy (anonymous):

Y=-2/3x+1 Y=-2/3*-6+1 Y=5 Y=-2/3*-3+1 Y=3 Y=-2/3*3+1 Y=-1 (-6,5)(-3,3)(3,-1)

OpenStudy (australopithecus):

for x = -6 y = 2/(3(-6) + 1) y = 2/((-18) + 1) y = 2/-17 y = -2/17

OpenStudy (australopithecus):

wait is the question \[y = \frac{-2}{3x+1} \] or \[ y = \frac{-2}{3x} + 1\]

OpenStudy (australopithecus):

and I made a mistake my answer should have been y = 2/17

OpenStudy (australopithecus):

if it is the first case

OpenStudy (australopithecus):

I guess I should have assumed it was the second case since there is a lack of brackets

OpenStudy (australopithecus):

y = -2/-18 + 1 y = 2/18 + 18/18 y = 20/18 y = 10/9

OpenStudy (australopithecus):

that is the solution for the second way your question can be interpreted, I dont know how you got 5 as a solution for x=-6

OpenStudy (cruffo):

Recall that the question mentioned that it was a linear equation :)

OpenStudy (australopithecus):

blah my reading comprehension is shot so burned out by school :) so it is really \[y = \frac{-2}{3}x + 1\]

OpenStudy (australopithecus):

y = -2*-6/3 + 1 y = 12/3 + 1 y = 5 alright yeah you are right good job :)

OpenStudy (anonymous):

thank you, i was starting to get a little confused lol. sometimes the numbers start running together in my head so i like to get on here and check my work

OpenStudy (australopithecus):

try https://www.wolframalpha.com

OpenStudy (australopithecus):

it is a great site for checking your work

OpenStudy (anonymous):

thank you